# DDAD: A Two-pronged Adversarial Defense Based on Distributional Discrepancy

### v1

@@ Abstract
Statistical adversarial data detection (SADD) detects whether an upcoming batch contains adversarial examples (AEs) by measuring the distributional discrepancies between clean examples (CEs) and AEs. In this paper, we reveal the potential strength of SADD-based methods by theoretically showing that minimizing distributional discrepancy can help reduce the expected loss on AEs. Nevertheless, despite these advantages, SADD-based methods have a potential limitation: they discard inputs that are detected as AEs, leading to the loss of clean information within those inputs. To address this limitation, we propose a two-pronged adversarial defense method, named Distributional-Discrepancy-based Adversarial Defense (DDAD). In the training phase, DDAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT, and then trains a denoiser by minimizing the MMD-OPT between CEs and AEs. In the inference phase, DDAD first leverages MMD-OPT to differentiate CEs and AEs, and then applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the distributional-discrepancy-based denoiser. Extensive experiments show that DDAD outperforms current state-of-the-art (SOTA) defense methods by notably improving clean and robust accuracy on CIFAR-10 and ImageNet-1K against adaptive white-box attacks.
@@ 1 Introduction
The discovery of adversarial examples (AEs) has raised a security concern for artificial intelligence techniques in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean examples (CEs), which can easily mislead a well-trained deep learning model to make wrong predictions. Considering the extensive use of deep learning systems, AEs pose a significant security threat for real-world applications (Sharif et al., 2016; Dong et al., 2019; Finlayson et al., 2019; Cao et al., 2021; Jing et al., 2021). Therefore, it is imperative to develop advanced defense methods to defend against AEs (Goodfellow et al., 2015; Madry et al., 2018; Zhang et al., 2019; Wang et al., 2020; Yoon et al., 2021; Nie et al., 2022).
Recently, statistical adversarial data detection (SADD) has gained increasing attention due to its effectiveness in detecting AEs (Gao et al., 2021; Zhang et al., 2023). Unlike other detection-based methods that train a detector for specific classifiers (Stutz et al., 2020; Deng et al., 2021; Pang et al., 2022b), SADD leverages statistical methods, for example, maximum mean discrepancy (MMD) (Gretton et al., 2012), to measure the discrepancies between the clean and adversarial distributions. Given the fact that clean and adversarial data are from different distributions, SADD-based methods have been shown empirically effective against adversarial attacks (Gao et al., 2021; Zhang et al., 2023).
@@
However, despite their effectiveness from both empirical and theoretical perspectives, detection-based methods (e.g., SADD-based methods) have a potential limitation: they discard inputs if they are detected as AEs, leading to the loss of clean information (e.g., semantic information) within those inputs. This issue is more prominent in SADD-based methods, where inputs are often processed in batches, potentially resulting in the unintended loss of some CEs along with AEs if a batch contains a mixture of CEs and AEs (Gao et al., 2021; Zhang et al., 2023). Furthermore, in many domains, obtaining large quantities of high-quality data is challenging due to factors such as cost, privacy concerns, or the rarity of specific data, for example, obtaining medical images for rare diseases is challenging (Litjens et al., 2017). As a result, all possible samples with clean information are critical in these data-scarce domains (Gandhar et al., 2024). Therefore, given the effectiveness of SADD-based methods, the above-mentioned challenges naturally lead us to pose the following question: Can we design an adversarial defense method that leverages the effectiveness of SADD-based methods, while at the same time, preserves all the data before feeding them into a classifier?
@@
The answer to this question is affirmative. Motivated by our theoretical analysis, we propose a two-pronged adversarial defense called Distributional-Discrepancy-based Adversarial Defense (DDAD). Specifically, we leverage an advanced MMD statistic (named MMD-OPT) in our pipeline, which is obtained by maximizing the testing power of MMD (see Algorithm 1). MMD-OPT serves two roles: in the training phase of the denoiser (see Algorithm 2), MMD-OPT serves as a ‘guider’ that can help minimize the distributional discrepancies between AEs and CEs. Then, by simultaneously minimizing the cross-entropy loss, we aim to train a denoiser that can minimize the distributional discrepancy towards the direction of making the classification correct; in the inference phase (see Section 4.3), MMD-OPT serves as a ‘detector’ that can help differentiate CEs and AEs. Then, our method applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the denoiser through distributional discrepancy minimization. We provide a visual illustration in Figure 1.
Through extensive evaluations on benchmark image datasets such as CIFAR-10 (Krizhevsky et al., 2009) and Imagenet-1K (Deng et al., 2009), we demonstrate the effectiveness of DDAD in Section 5. Compared to current state-of-the-art (SOTA) adversarial defense methods, DDAD can improve clean and robust accuracy by a notable margin against well-designed adaptive white-box attacks (see Section 5.2). Furthermore, experiments show that DDAD can generalize well against unseen transfer attacks (see Section 5.3).
@@
The success of DDAD in adversarial classification takes root in the following aspects: (1) minimizing distributional discrepancies has the potential to reduce the expected loss on AEs; (2) the two-pronged process combines the strengths of SADD-based and denoiser-based methods while also addressing their potential limitations: SADD-based methods can effectively distinguish AEs from CEs but discard the clean information within AEs. In contrast, denoiser-based methods can handle both data without re-training the downstream task model. However, they cannot distinguish AEs and CEs beforehand, which often results in a drop in clean accuracy. Our method, on the other hand, separates CEs and AEs in the inference phase, thereby keeping the accuracy for CEs nearly unaffected. At the same time, AEs can be properly handled by the denoiser; (3) compared to most denoiser-based methods that rely on density estimation (e.g., Nie et al. (2022) and Lee & Kim (2023)), learning distributional discrepancies is a simpler and more feasible task, especially on large-scale datasets such as ImageNet-1K.
2 Problem Setting
In this section, we discuss the problem setting for the adversarial classification in detail.
@@ 2 Problem Setting
We formalize our problem for $K$-class classification as follows. We define a domain as a pair consisting of a distribution $\mathcal{D}$ on inputs $\mathcal{X}$ and a labelling function $f:\mathcal{X}\to\{0,1,...,K\}$. Specifically, we consider a clean domain and an adversarial domain. The clean domain is denoted by $\langle\mathcal{D_{C}},f_{\mathcal{C}}\rangle$, and the adversarial domain is denoted by $\langle\mathcal{D_{A}},f_{\mathcal{A}}\rangle$. We define a hypothesis as a function $h:\mathcal{X}\rightarrow\{0,1,...,K\}$ from the hypothesis space $\mathcal{H}$. The probability according to the distribution $\mathcal{D}$ that a hypothesis $h$ disagrees with a labelling function $f$ (which can also be a hypothesis) is the risk:
@@
where $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))$ is a loss function that measures the disagreement between $h(\mathbf{x})$ and $f(\mathbf{x})$.
We consider the clean risk of a hypothesis as $R(h,f_{\mathcal{C}},\mathcal{D_{C}})$, and the adversarial risk as $R(h,f_{\mathcal{A}},\mathcal{D_{A}})$. In our problem, adversarial data are generated based on the given clean data. Therefore, $\mathcal{D_{C}}$ is fixed and we use $\mathbb{D}$ to represent a set of valid adversarial distributions such that all possible $\mathcal{D_{A}}\in\mathbb{D}$.
@@
3 Motivation from Theoretical Justification
In this section, we study a toy setting on the relationship between adversarial risk and distributional discrepancy, aiming to shed some light on designing effective adversarial defense methods.
Simplified problem setting. For simplicity, we analyze our problem for binary classification, i.e., a labelling function $f$ is simplified to $f:\mathcal{X}\to\{0,1\}$ and a hypothesis $h\in\mathcal{H}$ is simplified to $h:\mathcal{X}\rightarrow\{0,1\}$. The loss function is simplified to 0-1 loss (i.e., $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))=|h(\mathbf{x})-f(\mathbf{x})|$). Otherwise, other settings (e.g., the definition of risks) are the same as described in Section 2.
@@ 3 Motivation from Theoretical Justification
The proof of Theorem 1 can be found in Appendix A.
@@
Significance of distributional discrepancy to adversarial defense. In our problem, we use a practical setting that an attacker aims to attack a well-trained classifier on clean data (i.e., ideally the clean risk is minimized). According to Theorem 1, we have:
@@
Since $h_{\mathcal{C}}^{*}$, $f_{\mathcal{C}}$ and $\mathcal{D_{C}}$ are fixed, $R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})$ is possibly a small constant (according to Definition 2). In our problem, the objective of an attacker can be considered as finding an optimal $\mathcal{D_{A}}\in\mathbb{D}$ that maximizes $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$. Now, assume we have a detector that leverages the distributional discrepancies to identify AEs. Then, to break the defense, the attacker must generate AEs that could minimize the distributional discrepancies between CEs and AEs (i.e., to mislead the detector to identify AEs as CEs). However, according to Eq. 1, reducing the distributional discrepancy $d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})$ can help reduce adversarial risk $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$, which violates the objective of adversarial attacks. This intriguing phenomenon helps explain why SADD-based methods are effective against adaptive attacks in practice and inspires the design of our proposed method in this paper (see Section 4).
@@
Comparison with previous studies. Previous studies have attempted to use distributional discrepancy in adversarial defense. For example, at the early stage of AT, Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem. However, to the best of our knowledge, the relationship between adversarial risk and distributional discrepancy has not been well investigated yet from a theoretical perspective. In previous domain adaptation literature, the upper bound of the risk on the target domain is always bounded by one extra constant (Mansour et al., 2009; Ben-David et al., 2010), e.g., $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})\leq R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})+C$. This constant $C$ may prevent decreasing the risk on the target domain from minimizing the distributional discrepancy between the source domain and the target domain. By contrast, we treat adversarial classification as a special domain adaptation problem where the ground truth labelling functions are equivalent for both source and target domain. Based on this, we derive an upper bound without any extra constant, i.e., distributional discrepancy minimization can help reduce the expected loss on adversarial domain.
4 Methodology
Figure 1: The illustration of Distributional-Discrepancy-based Adversarial Defense (DDAD). In the training phase, DDAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT and then trains a denoiser by minimizing the MMD-OPT between CEs and AEs. Then, by simultaneously minimizing the cross-entropy loss, we aim to obtain a denoiser that can minimize the distributional discrepancy towards the direction of making the classification correct. In the inference phase, DDAD uses MMD-OPT to detect AEs and then denoises them instead of discarding them. Conversely, our method will directly feed detected CEs into the classifier.
@@ 4 Methodology
Motivated by our theoretical analysis in Section 3, we propose a two-pronged adversarial defense method called Distributional-Discrepancy-based Adversarial Defense (DDAD). In this section, we will first introduce the concepts of maximum mean discrepancy (MMD). This will be followed by a detailed discussion of the training and inference process of DDAD. We provide a visual illustration for DDAD in Figure 1 and a detailed description of mathematical notations in Appendix B.
@@ 4.1 Preliminary
Maximum mean discrepancy. The problem of using $L^{1}$ divergence in practice is that it does not any unbiased estimators. This is because the supremum can hardly be approximated by finite samples. Hence, in practice, it is challenging to estimate $L^{1}$ divergence accurately, especially in high-dimensional settings, where the bias and variance of the estimation can become significant. Therefore, in this paper, we use MMD to measure the distributional discrepancies between AEs and CEs. MMD has unbiased estimators and can effectively distinguish the difference between two distributions using small batches of data (Liu et al., 2020; Gao et al., 2021; Zhang et al., 2023). Following Gretton et al. (2012), let $\mathcal{X}\subset\mathbb{R}^{d}$ denote a separable metric space, and let $\mathbb{P}$ and $\mathbb{Q}$ represent Borel probability measures defined on $\mathcal{X}$. Given two sets of IID observations $S_{X}=\{\mathbf{x}^{(i)}\}_{i=1}^{n}$ and $S_{Z}=\{\mathbf{z}^{(i)}\}_{i=1}^{m}$ sampled from distributions $\mathbb{P}$ and $\mathbb{Q}$, respectively, kernel-based MMD (Borgwardt et al., 2006) quantifies the discrepancy between these two distributions:
$\textsc{MMD}(\mathbb{P},\mathbb{Q};\mathbb{H}_{k})$
$=\|\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}\|_{\mathbb{H}_{k}}$ (2)
@@
4.2 Training Process of DDAD
In this section, we discuss the training process of DDAD, which includes optimizing MMD and training the denoiser. We provide detailed algorithmic descriptions for the training process of DDAD in Algorithms 1 and 2.
Optimizing MMD. Following Liu et al. (2020), the test power of MMD can be maximized by maximizing the following objective (i.e., optimize $k_{\omega}$):

### v2

# One Stone, Two Birds: Enhancing Adversarial Defense Through the Lens of Distributional Discrepancy

@@ Abstract
Statistical adversarial data detection (SADD) detects whether an upcoming batch contains adversarial examples (AEs) by measuring the distributional discrepancies between clean examples (CEs) and AEs. In this paper, we explore the strength of SADD-based methods by theoretically showing that minimizing distributional discrepancy can help reduce the expected loss on AEs. Despite these advantages, SADD-based methods have a potential limitation: they discard inputs that are detected as AEs, leading to the loss of useful information within those inputs. To address this limitation, we propose a two-pronged adversarial defense method, named Distributional-discrepancy-based Adversarial Defense (DAD). In the training phase, DAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT, which is a stone that kills two birds. MMD-OPT first serves as a guiding signal to minimize the distributional discrepancy between CEs and AEs to train a denoiser. Then, it serves as a discriminator to differentiate CEs and AEs during inference. Overall, in the inference stage, DAD consists of a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the distributional-discrepancy-based denoiser. Extensive experiments show that DAD outperforms current state-of-the-art (SOTA) defense methods by simultaneously improving clean and robust accuracy on CIFAR-10 and ImageNet-1K against adaptive white-box attacks. Codes are publicly available at: https://github.com/tmlr-group/DAD.
@@ 1 Introduction
The discovery of adversarial examples (AEs) has raised a security concern for deep learning techniques in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean examples (CEs), which can easily mislead a well-trained deep learning model to make wrong predictions. Considering the extensive use of deep learning-based systems, AEs pose a significant security threat for real-world applications (Dong et al., 2019; Finlayson et al., 2019; Cao et al., 2021). Therefore, it is imperative to develop advanced defense methods to defend against AEs (Goodfellow et al., 2015; Madry et al., 2018; Zhang et al., 2019; Wang et al., 2020; Yoon et al., 2021; Nie et al., 2022).
Recently, statistical adversarial data detection (SADD) has gained increasing attention due to its effectiveness in detecting AEs (Gao et al., 2021; Zhang et al., 2023). Unlike other detection-based methods that train a detector for specific classifiers (Stutz et al., 2020; Deng et al., 2021; Pang et al., 2022b), SADD leverages statistical methods, for example, maximum mean discrepancy (MMD) (Gretton et al., 2012), to measure the discrepancies between clean and adversarial distributions. Given the fact that clean and adversarial data are from different distributions, SADD-based methods have been shown empirically effective against adversarial attacks (Gao et al., 2021; Zhang et al., 2023).
@@
However, despite their effectiveness from both empirical and theoretical perspectives, detection-based methods (e.g., SADD-based methods) have a potential limitation: they discard inputs if they are detected as AEs, leading to the loss of, e.g., semantic information within those inputs. This issue is more prominent in SADD-based methods, where inputs are often processed in batches, potentially resulting in the unintended loss of some CEs along with AEs if a batch contains a mixture of CEs and AEs (Gao et al., 2021; Zhang et al., 2023). Furthermore, in many domains, obtaining large quantities of high-quality data is challenging due to factors such as cost, privacy concerns, or the rarity of specific data, for example, obtaining medical images for rare diseases is challenging (Litjens et al., 2017). As a result, all possible samples with useful information are critical in these data-scarce domains (Gandhar et al., 2024). Therefore, given the effectiveness of SADD-based methods, the above-mentioned challenges naturally lead us to pose the following question: Can we design an adversarial defense method that leverages the effectiveness of SADD-based methods, while at the same time, preserves all the data before feeding them into a classifier?
@@
The answer to this question is affirmative. Motivated by our theoretical analysis, we propose a two-pronged adversarial defense called Distributional-discrepancy-based Adversarial Defense (DAD). Specifically, we leverage an advanced MMD statistic (named MMD-OPT) in our pipeline, which is obtained by maximizing the testing power of MMD (see Algorithm 1). MMD-OPT, as one stone, essentially kills two birds, i.e., serving two roles in DAD: in the training phase of the denoiser (see Algorithm 2), it acts as a guiding signal to minimize the distributional discrepancy between AEs and CEs. Then, by simultaneously minimizing the cross-entropy loss, we aim to train a denoiser that can minimize the distributional discrepancy towards the direction of making the classification correct. In the inference phase (see Section 4.3), MMD-OPT acts as a discriminator to distinguish between CEs and AEs. Then, our method applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the well-trained denoiser. We provide a visual illustration in Figure 1.
The success of DAD in adversarial classification takes root in the following aspects:
  • Statistical principle. Minimizing distributional discrepancies has been proven significant in controlling the expected loss on AEs. Thus, our new defense is built upon a solid theoretical foundation.
@@
  • One stone, two birds. DAD combines the strengths of SADD-based and denoiser-based methods while addressing their limitations: SADD-based methods will discard the useful information within AEs. In contrast, denoiser-based methods cannot distinguish AEs and CEs beforehand, which often results in a drop in clean accuracy. Our method, on the other hand, separates CEs and AEs in the inference phase, thereby keeping the accuracy for CEs nearly unaffected. At the same time, AEs can be properly handled by the denoiser.
  • Discrimination is easier than data generation. Compared to the most successful denoiser-based methods (known as adversarial purification) that rely on density estimation (e.g., Nie et al. (2022) and Lee & Kim (2023)), learning distributional discrepancies between AEs and CEs is a simpler and more feasible task, especially on large-scale datasets such as ImageNet-1K (Section 5.1).
Through extensive evaluations on benchmark image datasets such as CIFAR-10 (Krizhevsky et al., 2009) and Imagenet-1K (Deng et al., 2009), we demonstrate the effectiveness of DAD in Section 5. Compared to current state-of-the-art (SOTA) adversarial defense frameworks (i.e., adversarial training and adversarial purification), DAD can notably improve clean and robust accuracy simultaneously against well-designed adaptive white-box attacks (see Section 5.1). Furthermore, experiments show that DAD can generalize well against unseen transfer attacks (see Section 5.2).
2 Problem Setting and Assumptions
@@
In this section, we discuss the problem setting for the adversarial classification in detail. We formalize our problem for $K$-class classification as follows. We define a domain as a pair consisting of a distribution $\mathcal{D}$ on inputs $\mathcal{X}$ and a labelling function $f:\mathcal{X}\to\{1,...,K\}$. Specifically, we consider a clean domain and an adversarial domain. The clean domain is denoted by $\langle\mathcal{D_{C}},f_{\mathcal{C}}\rangle$, and the adversarial domain is denoted by $\langle\mathcal{D_{A}},f_{\mathcal{A}}\rangle$. We define a hypothesis as a function $h:\mathcal{X}\rightarrow\{1,...,K\}$ from the hypothesis space $\mathcal{H}$. The probability according to the distribution $\mathcal{D}$ that a hypothesis $h$ disagrees with a labelling function $f$ (which can also be a hypothesis) is the risk:
@@
where $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))$ is a loss function that measures the disagreement between $h(\mathbf{x})$ and $f(\mathbf{x})$. We consider the clean risk of a hypothesis as $R(h,f_{\mathcal{C}},\mathcal{D_{C}})$, and the adversarial risk as $R(h,f_{\mathcal{A}},\mathcal{D_{A}})$. In our problem, adversarial data are generated based on the given clean data. Therefore, $\mathcal{D_{C}}$ is fixed and we use $\mathbb{D}$ to represent a set of valid adversarial distributions such that all possible $\mathcal{D_{A}}\in\mathbb{D}$.
@@
3 Motivation from Theoretical Justification
In this section, we study the relationship between adversarial risk and distributional discrepancy, aiming to shed some light on designing effective adversarial defense methods.
3.1 Relation between labelling functions
We first reveal a simple yet important relation between labelling functions of CEs and AEs.
@@ 3.1 Relation between labelling functions
3.2 Theoretical justifications
For simplicity, we analyze our problem for binary classification, i.e., a labelling function $f$ is simplified to $f:\mathcal{X}\to\{0,1\}$ and a hypothesis $h\in\mathcal{H}$ is simplified to $h:\mathcal{X}\rightarrow\{0,1\}$. The loss function is simplified to 0-1 loss (i.e., $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))=|h(\mathbf{x})-f(\mathbf{x})|$). Otherwise, other settings (e.g., the definition of risks) are the same as described in Section 2.
@@ 3.2 Theoretical justifications
The proof of Theorem 1 can be found in Appendix A. Based on this theorem, we can easily find that distributional discrepancy is very important in adversarial defense.
Significance of distributional discrepancy to adversarial defense. We first give the definition of the well-trained classifier $h_{\mathcal{C}}^{*}$ in the adversarial attack scenarios.
@@
Normally, because an attacker aims to attack the well-trained classifier on clean data (i.e., ideally the clean risk is minimized), according to Theorem 1, we have
@@
Since $h_{\mathcal{C}}^{*}$, $f_{\mathcal{C}}$ and $\mathcal{D_{C}}$ are fixed, $R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})$ is possibly a small constant (according to Definition 2). In our problem, the objective of an attacker can be considered as finding an optimal $\mathcal{D_{A}}\in\mathbb{D}$ that maximizes $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$. Now, assume we have a detector that leverages the distributional discrepancies to identify AEs. Then, to break the defense, the attacker must generate AEs that could minimize the distributional discrepancies between CEs and AEs (i.e., to mislead the detector to identify AEs as CEs). However, according to Eq. (1), reducing the distributional discrepancy $d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})$ can help reduce adversarial risk $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$, which violates the objective of adversarial attacks. This intriguing phenomenon helps explain why SADD-based methods are effective against adaptive attacks in practice and inspires the design of our proposed method in this paper (see Section 4).
@@
Comparison with previous studies. Previous studies have attempted to use distributional discrepancy in adversarial defense. For example, at the early stage of adversarial training (AT), Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem. However, to the best of our knowledge, the relationship between adversarial risk and distributional discrepancy has not been well investigated yet from a theoretical perspective. In previous domain adaptation literature, the upper bound of the risk on the target domain is always bounded by one extra constant (Mansour et al., 2009; Ben-David et al., 2010), e.g., $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})\leq R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})+C$. This constant $C$ may prevent decreasing the risk on the target domain from minimizing the distributional discrepancy between the source domain and the target domain. By contrast, we find that adversarial classification is a special domain adaptation problem where the ground truth labelling functions are equivalent for both source and target domain. Based on this, we derive an upper bound without any extra constant, i.e., distributional discrepancy minimization can directly reduce the expected loss on adversarial domain.
4 A New Framework: Distributional Discrepancy-based Adversarial Defense
Figure 1: The illustration of Distributional-discrepancy-based Adversarial Defense (DAD). DAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT, which is a stone that kills two birds. MMD-OPT first serves as a guiding signal to minimize the distributional discrepancy between CEs and AEs to train a denoiser. In the inference stage, it also serves as a discriminator to differentiate CEs and AEs during inference. Then, our method applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the well-trained denoiser.
@@ 4 A New Framework: Distributional Discrepancy-based Adversarial Defense
Motivated by our theoretical analysis above, we propose a two-pronged adversarial defense framework called Distributional discrepancy-based Adversarial Defense (DAD) in this section. We first introduce the concepts of maximum mean discrepancy (MMD). This will be followed by a detailed discussion of the training and inference process of DAD which is illustrated in Figure 1. Detailed description of mathematical notations are in Appendix B.
@@ 4.1 Preliminary
Maximum mean discrepancy. The problem of using $L^{1}$ divergence in practice is that it does not have unbiased estimators. This is because the supremum can hardly be approximated by finite samples. Hence, in practice, it is challenging to estimate $L^{1}$ divergence accurately, especially in high-dimensional settings, where the bias and variance of the estimation can become significant. Therefore, in this paper, we use MMD to measure the distributional discrepancies between AEs and CEs. MMD has unbiased estimators and can effectively distinguish the difference between two distributions using small batches of data (Liu et al., 2020, 2021a; Zhang et al., 2023). Following Gretton et al. (2012), let $\mathcal{X}\subset\mathbb{R}^{d}$ denote a separable metric space, and let $\mathbb{P}$ and $\mathbb{Q}$ represent Borel probability measures defined on $\mathcal{X}$. Given two sets of IID observations $S_{X}=\{\mathbf{x}^{(i)}\}_{i=1}^{n}$ and $S_{Z}=\{\mathbf{z}^{(i)}\}_{i=1}^{m}$ sampled from distributions $\mathbb{P}$ and $\mathbb{Q}$, respectively, kernel-based MMD (Borgwardt et al., 2006) quantifies the discrepancy between these two distributions:
$\textsc{MMD}(\mathbb{P},\mathbb{Q};\mathbb{H}_{k})=\|\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}\|_{\mathbb{H}_{k}}$
@@
3:  for epoch $=1,...,T$ do
6:    $k_{\omega}\leftarrow$ kernel function with parameters $\omega$ using Eq. (3);
7:    $M(\omega)\leftarrow\widehat{\textsc{MMD}}^{2}_{\rm u}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. (2);
8:    $V_{\lambda}(\omega)\leftarrow\hat{\sigma}_{\lambda}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. (5);
9:    $\hat{J}_{\lambda}(\omega)\leftarrow M(\omega)/\sqrt{V_{\lambda}(\omega)}$ using Eq. (4.2);
11:  end for
12:  Output: $k^{*}_{\omega}$
Algorithm 1 Optimizing MMD (Liu et al., 2020).
4.2 Training Process of DAD
In this section, we discuss the training process of DAD, which includes optimizing MMD and training the denoiser. We provide detailed algorithmic descriptions for the training process of DAD in Algorithms 1 and 2.
One stone: optimized MMD. Following Liu et al. (2020), the test power of MMD can be maximized by maximizing the following objective (i.e., optimize $k_{\omega}$):